OpenStudy (anonymous):

Find the average velocity vector during the time interval between t=2 and t=3 v(t)= 2i - 6tj

OpenStudy (anonymous):

I've substituted the values 2 and 3 into the velocity vector and got the answers; V(2) = 2i - 12j v(3) = 21 -18j

OpenStudy (anonymous):

then velocity average = [v(3) -V(2)]/2

OpenStudy (anonymous):

but my answers doesn't match the back of the book;(

OpenStudy (anonymous):

the average velocity from t = a to t = b is:\[\frac{v(b)-v(a)}{b-a}\]is this the formula you are using?

OpenStudy (anonymous):

oh i thought you just divide v(b)-v(a) by 2

OpenStudy (anonymous):

you need to divide it by the amount of time that elapsed during your interval, which is 3-2=1 second. i might be wrong, im not much of a physics person.

OpenStudy (anonymous):

Guys, you can't average at the beginning and the end because velocity is not increasing linearly. You need to integrate, evaluate the definite integral and then divide by the total time.

OpenStudy (anonymous):

my answer was -15j but the back of the book says 2i-15j

OpenStudy (aravindg):

pls giv medal to me

OpenStudy (anonymous):

ah i see, my bad >.<

OpenStudy (anonymous):

oh wait, sorry, it does say t and not t^2, so yeah, averaging the endpoints works. my bad

OpenStudy (anonymous):

lol mi sorta confused

OpenStudy (anonymous):

is my method correct?

OpenStudy (anonymous):

yes, you can evaluate the vector at each endpoint, add them together, and then divide by the time difference

OpenStudy (anonymous):

if the answer is 2i-15j, it seems you need to integrate, in which you get 2ti-3t^2j, then plug in the values t = 2 and 3. This gives 4i-12j 6i-27j then subtract the bottom from the first and divide by 3-2=1. that gives 2i-15j

OpenStudy (anonymous):

Formulas \[Average Velocity= \frac{Final Position- Initial Position}{Time Interval}\] So \[Postition =\int Velocity \;\;dt\] \[Position=\int (2)i-(6t)j \;\;dt \;\;= (2t)i-(3t^{3})j\] So, \[Average \; velocity= \frac{\int_{2}^{3} [(2)i-(6t)j]dt}{3-2} \;\;dt \] \[Average \; velocity= (2t)i-(3t^{3})j \|^{3}_{2}\] \[Average \; velocity=(2)i-(3(3^2-2^2))j= 2i-15j\] and Bingo.

OpenStudy (anonymous):

a much better explanation lol

OpenStudy (anonymous):

wow thanks heaps guys:)